a gas mixture has following composition by volume:
          SO2-8.5%
           O2- 10%
           N2-81.5%
calculate:
  1. composition by weight
  2. average molecular weight of mixture
  3. density of gas mixture at a temparature 200 degree celcius and 202 kilo pascal

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Apply ideal gas eqn. for each gas,
For SO2,
2*8.5 = n1*R*473
n1= 0.4337 and weight1 = 27.75 gm
Similaly n2 = 20/R*473 and weight2 = 16.48 gm
n2 = 0 .515 moles
and
n3 = 4.197
and weight3 = 117.515 gm
 
Avg. molecular weight = (117.515 + 16.48 + 27.75)/3 = > 53.91
and
density = PM/RT = > 2*53.91/0.0821*473 => 2.77 g/l